# Robust Adaptive Safe Robotic Grasping with Tactile Sensing

### v1

@@ Abstract
Robotic grasping requires safe force interaction to prevent a grasped object from being damaged or slipping out of the hand. In this vein, this paper proposes an integrated framework for grasping with formal safety guarantees based on Control Barrier Functions. We first design contact force and force closure constraints, which are enforced by a safety filter to accomplish safe grasping with finger force control. For sensory feedback, we develop a technique to estimate contact point, force, and torque from tactile sensors at each finger. We verify the framework with various safety filters in a numerical simulation under a two-finger grasping scenario. We then experimentally validate the framework by grasping multiple objects, including fragile lab glassware, in a real robotic setup, showing that safe grasping can be successfully achieved in the real world. We evaluate the performance of each safety filter in the context of safety violation and conservatism, and find that disturbance observer-based control barrier functions provide superior performance for safety guarantees with minimum conservatism. The demonstration video is available at https://youtu.be/Cuj47mkXRdg.
@@ I Introduction
The human hand is capable of adapting to a wide range of complex objects and performing different tasks in daily life robustly [1], since it allows smooth and safe force interaction [2]. Inspired by this, robotic grasping has been investigated across many human-oriented applications such as industrial assembly [3], packing of groceries [4], and household tasks [5]. Despite the significant progress of robotic grasping over the past decades, the grasping research still revolves around determining proper finger postures for grasping from a kinematic perspective [6][7]. At the same time, contact dynamics are often overlooked in existing grasping methods, which are not sufficient for achieving dexterous grasping [8]. This limitation prevents robotic grasping from being applied to more general tasks that require safe behaviors, as they cannot prevent slippage between the hand and object.
To mitigate this limitation, some researchers have focused on force-based grasping techniques. In [9], they propose an adaptive force control framework that allows the hand to be compliant by mapping the human hand posture data into the desired force commands. To prevent undesired slippage, [10] proposes a robust force controller for grasping to maintain secure grasping by introducing force feedback. Similarly, [11] demonstrates a low-level impedance-based controller that incorporated task-based search strategies, comparing its effectiveness on a peg-in-hole task using various robotic hands. Furthermore, a method for regulating the grasping force based on tactile sensors is proposed in [12] to increase grasping stability for unknown objects by leveraging data-driven models like deep neural networks or Gaussian mixture models to detect and estimate contact status and force. Others have demonstrated that online adaptive control strategies can account for disturbances and model error [13] using dynamical systems-based [14] approaches for dexterous manipulation.
The modeling of contact dynamics has also played a key role in many robust grasping works. In [15], internal and friction forces of a grasped object are determined in order to provide stable finger gaits so that the proposed impedance controller can handle the dynamic changes in object state during reconfiguration of the fingers. Similarly, [16] proposes a hybrid position/force-controller used to generate a finger gaiting sequence while considering dynamic constraints imposed by friction cones, enabling stable control over object pose.
@@
Such force-based grasping methods have primarily focused only on stable grasping, neglecting safety considerations, which are crucial in real world applications. For instance, safety regulations are essential for laboratory automation [17] and space applications including robotic hands [18]. Therefore, ensuring safety in robotic grasping remains an open challenge beyond merely achieving compliant manipulation. To that end, a few works exist, proposing methods such as safe grasping with impact force measurements [19], safety-optimized strategies with visual depth data and false-positive detection for human safety [20], and an integrated framework for predicting safe grasping forces using transformers [21].
@@
Nevertheless, to the best of our knowledge, few studies take into account formal safety guarantees (e.g., force regulation and closure to avoid slippage) for force-based grasping control; thus, imbuing grasping algorithms with formal safety guarantees should be investigated for broader use cases such as medical, space, and chemical laboratory applications. Safety guarantees can be efficiently achieved by using reachability analysis [22], set invariance techniques [23], and control barrier functions (CBFs) [24][25][26]. In this paper, we focus on CBF-based controllers to ensure safety guarantees in robotic grasping. CBF-based controllers enforce the forward invariance of a defined safe set and can be efficiently implemented using quadratic programming, enabling their widespread use in many applications [27][28]. The use of CBFs for safe grasping was first proposed by [29]. Their methodology used robust CBFs in order to handle external disturbances, and consequently, achieved safe grasping while avoiding undesired slippage.
@@ I-A Contributions
  • We present a safe grasping framework with formal safety guarantees, based on CBFs schemes as safety filters. The scalability of the framework presented in this paper is demonstrated by applying several safety filters, including CBF [24], robust adaptive CBF [25] (RaCBF), robust CBF (RCBF) [26], and disturbance observer CBF (DOBCBF) [30].
@@
We show significant improvements in our framework compared to [29] with respect to adaptation, conservatism, and grasping force regulation. Our method incorporates adaptation to parametric uncertainty and while estimating external disturbances, making it less conservative than [29]. Additionally, since [29] maintains a strong grasping force to prevent slipping without regulating the force, it may lead to applying excessive force and potentially damaging the grasped object. In this work, constraints including contact force and force closure are simultaneously considered.
This paper is organized as follows. Section II defines a problem solved in the paper and Section III provides safety filters and a disturbance observer as preliminaries. Section IV presents the proposed framework. Subsequently, we verify our framework in numerical simulations in Section V and validate it in the real multiple experiments in Section VI. Lastly, we conclude our paper with future works in Section VII.
@@ II Problem Formulation
where $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state, $f:\mathcal{X}\rightarrow\mathbb{R}^{n}$, $g_{1}:\mathcal{X}\rightarrow\mathbb{R}^{n\times m}$, and $g_{2}:\mathcal{X}\rightarrow\mathbb{R}^{n\times q}$ are locally Lipschitz continuous functions, $F:\mathcal{X}\rightarrow\mathbb{R}^{n\times k}$ is a smooth function with $F(\bm{0})=\bm{0}$, $\bm{\theta^{*}}\in\Theta\subset\mathbb{R}^{k}$ is the vector of constant unknown parameters, and $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represents control force input, and $\bm{d}\in\mathbb{R}^{q},||\bm{d}||\leq w_{0},||\dot{\bm{d}}||\leq w_{1}$ is a bounded external disturbance. We address the following problem:
@@ III-A Contact Dynamics Models
Consider the following second-order Kelvin-Voigt model [31] for contact forces:
@@
where $m_{o}\in\mathbb{R}_{>0}$ is mass, and $\bm{u}\in\mathbb{R}^{m}$ is the control force from each finger, and $\bm{I}_{m}\in\mathbb{R}^{m\times m}$ is an $m\times m$ dimensional identity matrix. For the system (2), we employ the following model:
@@
$g_{1}(\bm{x})$ $=\begin{bmatrix}\bm{0}\\ \frac{1}{m_{o}}\bm{I}_{m}\end{bmatrix},\quad g_{2}(\bm{x})=\begin{bmatrix}\bm{0}\\ \frac{1}{m_{o}}\bm{I}_{q}\end{bmatrix}.$ (5)
Lastly, the soft finger friction cone model is defined as follows [32]:
@@ III-B Robust Adaptive Control Barrier Functions
$\qquad\qquad\qquad\geq-\alpha\bigg{(}h_{r}(\bm{x},\bm{\hat{\theta}})-\frac{1}{2}\bm{\tilde{\vartheta}}^{\top}\Gamma^{-1}\bm{\tilde{\vartheta}}\bigg{)},$ (8)
with $\lambda(x,\bm{\theta})\triangleq{\bm{\theta}}-\Gamma\Big{(}\frac{\partial h_{r}}{\partial\bm{\theta}}(x,\bm{\theta})\Big{)}^{\top}$, for any $\bm{\hat{\theta}}\in\Theta$ satisfying the adaptation law $\dot{\bm{\hat{\theta}}}=-\Gamma\Big{(}\frac{\partial h_{r}}{\partial\bm{x}}(\bm{x},\bm{\hat{\theta}})F(\bm{x})\Big{)}^{\top}$, and where $\Gamma\in\mathbb{R}^{n\times n}$ is an adaptive gain whose minimum eigenvalue satisfies $\lambda_{\text{min}}(\Gamma)\geq\frac{||\bm{\tilde{\vartheta}}||^{2}}{2h_{r}(x,\bm{\theta})}$. The system (1) is safe with respect to (7) if $h_{r}$ is a RaCBF.
@@ III-D Nonlinear Disturbance Observer (NDOB)
Consider the uncertain system (2). Assuming that $||\bm{d}(t)||\leq w_{0}$ and $||\dot{\bm{d}}(t)||\leq w_{1}$, we deploy the following nonlinear disturbance observer proposed by [33]:
@@
where $\alpha_{d}$ is a constant positive adjustable parameter, and $P(\bm{x})$ is the gain function for the estimator satisfying $\frac{\partial P}{\partial\bm{x}}=L_{d}(\bm{x})$. $L_{d}(\bm{x})$ is the predefined gain function satisfying $\bm{e}_{d}^{\top}\bm{e}_{d}\leq\bm{e}_{d}^{\top}L_{d}(\bm{x})g_{2}(\bm{x})\bm{e}_{d}$ [34], where $\bm{e}_{d}=\bm{d}-\hat{\bm{d}}$ is the estimation error, and $\bm{e}_{d}$ is uniformly bounded by $||\bm{e}_{d}(t)||\leq\sqrt{\frac{2ck||\bm{e}_{d}(0)||^{2}e^{-2kt}+w^{2}_{1}(1-e^{-2kt})}{2ck}}:=M_{d}$, where $k=\alpha_{d}-\frac{c}{2}$, and $c$ is a non-negative parameter that satisfies $c<2\alpha_{d}$.
@@ IV-A1 Contact Force constraint
$\mathcal{H}_{f_{\textnormal{c}}}=\mathcal{H}_{f_{\textnormal{min}}}\times\mathcal{H}_{f_{\textnormal{max}}},$ (12)
@@
$\mathcal{H}_{f_{\textnormal{min}}}$ =$=\{\bm{x}\text{ }|\text{ }\bm{f}_{c}(\bm{x})-\bm{f}_{\textnormal{min}}\geq 0\},$ (13)
$\mathcal{H}_{f_{\textnormal{max}}}$ =$=\{\bm{x}\text{ }|\text{ }-\bm{f}_{c}(\bm{x})+\bm{f}_{\textnormal{max}}\geq 0\}.$ (14)
@@ IV-A2 Force Closure
To prevent slipping during grasping, we first consider a Coulomb friction model given in (6) and linearize it to be conservatively approximated in a set of tangential planes to the original cone as follows [35]:
@@
$\mathcal{H}=\mathcal{H}_{f_{\textnormal{c}}}\times\mathcal{H}_{s},$ (18)
@@ IV-B Safety Filters for Safe Grasping
$u_{\mathsf{safe}}=\quad\operatorname*{arg\,min}_{u\in\mathcal{U}}\frac{1}{2}{||u-u_{\mathsf{nominal}}||}^{2},$ (19)
where $u_{\mathsf{nominal}}\in\mathbb{R}^{m}$ is a nominal control input, and $u_{\mathsf{safe}}\in\mathbb{R}^{m}$ is a safe control input from each safety filter. The constraint (18) is easily enforced by a standard CBF [24] for the system (2) without considering $\bm{d}$. However, since standard CBFs are vulnerable to model uncertainties, we use following methods in our framework: RaCBFs [25], RCBFs [26], and DOBCBFs [30]. RaCBFs resolve inherent parametric model uncertainty for the system (4), and can provide a safe input by imposing $\mathcal{H}$ into (1) with $\bm{\tilde{\vartheta}}$. For more general use cases, RCBFs for the system (5) can attenuate bounded external disturbances by applying $\mathcal{H}$ into (9) with $w_{0}$. Lastly, the disturbance observer (10) can estimate the external disturbance in the system (2). $P(\bm{x})$ and $L_{d}$ are problem-specific gain functions that we choose to be:
@@
Subsequently, we include the estimated disturbance into safety condition for DOBCBFs [33] as follows:
@@
with $\psi_{0}=L_{f}h(\bm{x})+L_{g_{2}}h(\bm{x})\hat{\bm{d}}-\frac{w^{2}_{1}}{2c\beta}-\frac{\beta||L_{g_{2}}h||^{2}}{4\nu-2c-2\alpha}+\alpha(h(\bm{x}))$ where the positive constant $\beta,\nu$ such that $\nu>\frac{\alpha+c}{2}$ and $\beta>\frac{||\bm{e}_{d}(0)||^{2}}{2h(\bm{x}(0))}$. In the subsequent section, a fingertip force controller for the robotic hand is designed to track a safe input generated by each safety filter.
@@ IV-C Fingertip Contact Force Controller
[…]
@@ V Simulation
  • normal reaction forces are equal to or greater than $-6$N.
@@
Despite these positive results, we anticipate that determining the true friction coefficient in the real experiment is non-trivial unless we estimate it; thus, we assume that the friction coefficient error, $|\mu-\hat{\mu}|\leq e_{\textnormal{max}},$ $e_{\textnormal{max}}\geq 0$ is conservatively bounded as in [29] for the following real-world experiments.
@@ VI Hardware Experiments
In this section, we show the experimental results of our proposed method, demonstrating the safe grasping of various objects including a fragile object. To this end, we primarily use two fingers of the Shadow hand equipped with tactile sensors to grasp an object and control the front finger to regulate contact normal force. Lastly, we evaluate each safety filter in terms of safety violations and conservatism, and discuss the overall limitations of our framework.
@@
Safety Filters CBF [24] -0.1983 0.1624 -0.034 -0.159 0.448 -0.140 -0.085 1.928 -0.426 -0.353 0.257 -0.037
RaCBF [25] 0.049 0.214 0.018 0.195 0.641 0.0524 0.1871 2.898 0.544 0.01 0.3 0.005
RCBF [26] 0.248 0.208 0.012 0.252 0.665 0.077 0.195 2.852 0.498 0.064 0.307 0.013
DOBCBF [30] 0.007 0.196 0.001 0.046 0.64 0.0520 0.078 2.795 0.4401 0.003 0.305 0.011

### v2

@@ Abstract
Robotic grasping requires safe force interaction to prevent a grasped object from being damaged or slipping out of the hand. In this vein, this paper proposes an integrated framework for grasping with formal safety guarantees based on Control Barrier Functions. We first design contact force and force closure constraints, which are enforced by a safety filter to accomplish safe grasping with finger force control. For sensory feedback, we develop a technique to estimate contact point, force, and torque from tactile sensors at each finger. We verify the framework with various safety filters in a numerical simulation under a two-finger grasping scenario. We then experimentally validate the framework by grasping multiple objects, including fragile lab glassware, in a real robotic setup, showing that safe grasping can be successfully achieved in the real world. We evaluate the performance of each safety filter in the context of safety violation and conservatism, and find that disturbance observer-based control barrier functions provide superior performance for safety guarantees with minimum conservatism.
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
The human hand is capable of adapting to a wide range of complex objects and performing different tasks in daily life robustly [Liu2014], since it allows smooth and safe force interaction [Tong2024]. Inspired by this, robotic grasping has been investigated across many human-oriented applications such as industrial assembly [LiTME2019], packing of groceries [ChenRoboSoft2024], and household tasks [Jessica2024]. Despite the significant progress of robotic grasping over the past decades, the grasping research still revolves around determining proper finger postures for grasping from a kinematic perspective [Zhao2024ICRA][Albert2023IROS]. At the same time, contact dynamics are often overlooked in existing grasping methods, which are not sufficient for achieving dexterous grasping [TangICRA2024]. This limitation prevents robotic grasping from being applied to more general tasks that require safe behaviors, as they cannot prevent slippage between the hand and object.
To mitigate this limitation, some researchers have focused on force-based grasping techniques. In [ZengIROS2021], they propose an adaptive force control framework that allows the hand to be compliant by mapping the human hand posture data into the desired force commands. To prevent undesired slippage, [CostanzoICRA2020] proposes a robust force controller for grasping to maintain secure grasping by introducing force feedback. Similarly, [VanTRO2018] demonstrates a low-level impedance-based controller that incorporated task-based search strategies, comparing its effectiveness on a peg-in-hole task using various robotic hands. Furthermore, a method for regulating the grasping force based on tactile sensors is proposed in [Deng2020] to increase grasping stability for unknown objects by leveraging data-driven models like deep neural networks or Gaussian mixture models to detect and estimate contact status and force. Others have demonstrated that online adaptive control strategies can account for disturbances and model error [Khadivar2023TRO] using dynamical systems-based [Billard2011DS] approaches for dexterous manipulation.
The modeling of contact dynamics has also played a key role in many robust grasping works. In [Pfanne2020RAL], internal and friction forces of a grasped object are determined in order to provide stable finger gaits so that the proposed impedance controller can handle the dynamic changes in object state during reconfiguration of the fingers. Similarly, [Xue2023IROS] proposes a hybrid position/force-controller used to generate a finger gaiting sequence while considering dynamic constraints imposed by friction cones, enabling stable control over object pose.
@@
Such force-based grasping methods have primarily focused only on stable grasping, neglecting safety considerations, which are crucial in real-world applications. For instance, safety regulations are essential for laboratory automation [WOLF202218] and space applications including robotic hands [Hadi2024]. Therefore, ensuring safety in robotic grasping remains an open challenge beyond merely achieving compliant manipulation. To that end, a few works exist, proposing methods such as safe grasping with impact force measurements [Mavrakis2017IROS], safety-optimized strategies with visual depth data and false-positive detection for human safety [Li2024UR], and an integrated framework for predicting safe grasping forces using transformers [Han2024ASME].
@@
Nevertheless, to the best of our knowledge, few studies take into account formal safety guarantees (e.g., force regulation and closure to avoid slippage) for force-based grasping control; thus, imbuing grasping algorithms with formal safety guarantees should be investigated for broader use cases such as medical, space, and chemical laboratory applications. Safety guarantees can be efficiently achieved by using reachability analysis [Fisac2019], set invariance techniques [Wolff2004], and control barrier functions (CBFs) [Ames2019CBFtheoryandapplications][lopez2020robust][JANKOVIC2018359]. In this paper, we focus on CBF-based controllers to ensure safety guarantees in robotic grasping. CBF-based controllers enforce the forward invariance of a defined safe set and can be efficiently implemented using quadratic programming, enabling their widespread use in many applications [dawson2022barrier][YKECC2024]. The use of CBFs for safe grasping was first proposed by [Shaw2021]. Their methodology used robust CBFs in order to handle external disturbances, and consequently, achieved safe grasping while avoiding undesired slippage.
@@ I-A Contributions
  • We present a safe grasping framework with formal safety guarantees, based on CBFs schemes as safety filters. The scalability of the framework presented in this paper is demonstrated by applying several safety filters, including CBF [Ames2019CBFtheoryandapplications], robust adaptive CBF [lopez2020robust] (RaCBF), robust CBF (RCBF) [JANKOVIC2018359], and disturbance observer CBF (DOBCBF) [Wang2023].
@@
We show significant improvements in our framework compared to [Shaw2021] with respect to adaptation, conservatism, and grasping force regulation. The proposed framework incorporates adaptation to parametric uncertainty while estimating external disturbances, making it less conservative than [Shaw2021]. Additionally, since [Shaw2021] maintains a strong grasping force to prevent slipping without regulating the force, it may lead to applying excessive force and potentially damaging the grasped object. In this work, constraints including contact force and force closure are simultaneously considered.
This paper is organized as follows. Section II defines a problem solved in the paper and Section III provides safety filters and a disturbance observer as preliminaries. Section IV presents the proposed framework. Subsequently, we verify our framework in numerical simulations in Section V and validate it in the real multiple experiments in Section VI. Lastly, we conclude our paper in Section VII.
@@ II Problem Formulation
[…]
@@ III-A Contact Dynamics Models
Consider the following second-order Kelvin-Voigt model [kelvinModel] for contact forces:
@@
where $m_{o}\in\mathbb{R}_{>0}$ is mass, and $\bm{u}\in\mathbb{R}^{m}$ is the control force from each finger. $\bm{I}_{m}\in\mathbb{R}^{m\times m}$ represents an $m\times m$ dimensional identity matrix where $n=m$ in this case. For the system (2), we employ the following model:
@@
$g_{1}(\bm{x})$ $=\begin{bmatrix}\bm{0}\\ \frac{1}{m_{o}}\bm{I}_{m}\end{bmatrix},\quad g_{2}(\bm{x})=\begin{bmatrix}\bm{0}\\ \frac{1}{m_{o}}\bm{I}_{q}\end{bmatrix}$, (5)
where $\bm{I}_{q}\in\mathbb{R}^{n\times m}$ is a disturbance characteristic matrix.
Lastly, the soft finger friction cone model is defined as follows [Prattichizzo2016]:
@@ III-B Robust Adaptive Control Barrier Functions
[…]
@@ III-D Nonlinear Disturbance Observer (NDOB)
Consider the uncertain system (2). Assuming that $\bm{d}(t)||\leq w_{0}$ and $\dot{\bm{d}}(t)||\leq w_{1}$, we deploy the following nonlinear disturbance observer proposed by [DOBChens2004]:
@@
[…]
@@ IV-A1 Contact Force constraint
$\mathcal{H}_{f_{\textnormal{c}}}=\mathcal{H}_{f_{\textnormal{min}}}\cap\mathcal{H}_{f_{\textnormal{max}}},$ (12)
@@
$\mathcal{H}_{f_{\textnormal{min}}}$ $=\{\bm{x}\text{ }|\text{ }\bm{f}_{c}(\bm{x})-\bm{f}_{\textnormal{min}}\geq 0\},$ (13)
$\mathcal{H}_{f_{\textnormal{max}}}$ $=\{\bm{x}\text{ }|\text{ }-\bm{f}_{c}(\bm{x})+\bm{f}_{\textnormal{max}}\geq 0\}.$ (14)
@@ IV-A2 Force Closure
To prevent slipping during grasping, we first consider a Coulomb friction model given in (6) and linearize it to be conservatively approximated in a set of tangential planes to the original cone as follows [analysishandsJeffrey1986]:
@@
$\mathcal{H}=\mathcal{H}_{f_{\textnormal{c}}}\cap\mathcal{H}_{s},$ (18)
@@ IV-B Safety Filters for Safe Grasping
$\bm{u}_{\mathsf{safe}}=\quad\operatorname*{arg\,min}_{\bm{u}\in\mathcal{U}}\frac{1}{2}{||\bm{u}-\bm{u}_{\mathsf{nominal}}||}^{2},$ (19)
where $\bm{u}_{\mathsf{nominal}}\in\mathbb{R}^{m}$ is a nominal control input, and $\bm{u}._{\mathsf{safe}}\in\mathbb{R}^{m}$ is a safe control input from each safety filter. The constraint (18) is easily enforced by a standard CBF [Ames2019CBFtheoryandapplications] for the system (2) without considering $\bm{d}$. However, since standard CBFs are vulnerable to model uncertainties, we use following methods in our framework: RaCBFs [lopez2020robust], RCBFs [JANKOVIC2018359], and DOBCBFs [Wang2023]. RaCBFs resolve inherent parametric model uncertainty for the system (4), and can provide a safe input by imposing $\mathcal{H}$ into (1) with $\bm{\tilde{\vartheta}}$. For more general use cases, RCBFs for the system (5) can attenuate bounded external disturbances by applying $\mathcal{H}$ into (9) with $w_{0}$. Lastly, the disturbance observer (10) can estimate the external disturbance in the system (2). $P(\bm{x})$ and $L_{d}$ are problem-specific gain functions that we choose to be:
@@
Subsequently, we include the estimated disturbance into safety condition for DOBCBFs [DOBChens2004] as follows:
@@
[…]
@@ IV-C Fingertip Contact Force Controller
$\bm{q}_{\textnormal{ref}}=\bm{q}_{\textnormal{d}}+k_{\textrm{p}}\tilde{\bm{\tau}}+k_{\textrm{i}}\int\tilde{\bm{\tau}}dt+k_{\textrm{d}}\dot{\tilde{\bm{\tau}}}$, (24)
@@ V Simulation
  • normal reaction forces are equal to or greater than 6-6N.
@@
Despite these positive results, we anticipate that determining the true friction coefficient in the real experiment is non-trivial unless we estimate it; thus, we assume that the friction coefficient error, $|\mu-\hat{\mu}|\leq e_{\textnormal{max}},$ $e_{\textnormal{max}}\geq 0$ is conservatively bounded as in [Shaw2021] for the following real-world experiments.
@@ VI Hardware Experiments
In this section, we show the experimental results of the proposed method, demonstrating the safe grasping of various objects including a fragile object. To this end, we primarily use two fingers of the Shadow hand equipped with tactile sensors to grasp an object and control the front finger to regulate contact normal force. Lastly, we evaluate each safety filter in terms of safety violations and conservatism, and discuss the overall limitations of our framework. The demonstration video is available at the link [^1] https://youtu.be/Cuj47mkXRdg.
@@
Safety
Filters
CBF [Ames2019CBFtheoryandapplications] -0.1983 0.1624 -0.034 -0.159 0.448 -0.140 -0.085 1.928 -0.426 -0.353 0.257 -0.037
RaCBF [lopez2020robust] 0.049 0.214 0.018 0.195 0.641 0.0524 0.1871 2.898 0.544 0.01 0.3 0.005
RCBF [JANKOVIC2018359] 0.248 0.208 0.012 0.252 0.665 0.077 0.195 2.852 0.498 0.064 0.307 0.013
DOBCBF [Wang2023] 0.007 0.196 0.001 0.046 0.64 0.0520 0.078 2.795 0.4401 0.003 0.305 0.011